# Vibration-based damage detection of a trainer jet via multiple input tangential interpolation

### v3

@@ 2.3 A novel two-step modal & damage identification method
Noting that all SHM methods depend on identified modal parameters, an accurate and efficient SHM method can be developed by merging the previous two tools, the iLF and the MTMAC.
The full workflow for this novel SHM method is:
@@
    Modal identification of a given system;
@@
    The results of the above are used to validate the MTMAC for damage assessment and quantification;
@@
    For completeness, the COMAC is used to locate the damage.
In order to benchmark the above method, the more conventional LSCE and SSI-CVA methods are used, and the results are compared to iLF. Furthermore, traditional vibration-based damage assessment, quantification, and localisation comparing $\omega_{n}$ and $\bm{\phi}_{n}$ is used to validate the MTMAC as a damage index.
@@ 3 Numerical case study
To assess the suitability of the iLF and the MTMAC for damage detection, a numerical case study of a 2D hollow rectangular cross-section beam is introduced. The beam overall geometry cross-section is shown in Figure 1. It is made of Aluminium, with a Young’s modulus (E) of 69 GPa and density ($\rho$) of 2700 kgm-3, and a length (L) of 1.8 m. The beam is discretised, in MATLAB, into four 2D (y and z displacements) Euler-Bernoulli beam elements of equal length. Sample beam element mass and stiffness matrices are available in Appendix A.
Figure 1: Numerical case study: schematic of the 2D beam with the dimensions of the cross-section. Not in scale.
[…]
@@
The maximum error of the identified parameters ( and ) reported in Table 1 is 0%. However, LSCE fails to identify lower modes in higher-dimensional systems with closely spaced modes, such as modes 1-2, as well as modes 3-6. These results align with expectations, as this issue is well-documented in the literature [43, 3]. Conversely, Figure 2 clearly shows that the stabilisation diagrams from iLF (Figure 2(a)) and SSI-CVA (Figure 2(c)) can accurately identify the stable modes, while for LSCE (Figure 2(b)) this is troublesome, resulting in only ten (out of sixteen) modes identified as stable. It should be noted, although not included in Table 1, that the MAC between the analytical $\bm{\phi}_{n}$ and those identified via iLF and SSI-CVA is equal to, or very close to, 1. On the other hand, the LSCE identified $\bm{\phi}_{n}$ MAC, when compared to the analytical values, never exceeds 0.58.
@@ 3.1 Investigation of damage effects
Given the good accuracy and precision shown for the baseline case by the iLF when compared to the SSI-CVA and LSCE methods, only the iLF-derived modal parameters are used for damage assessment in the numerical case study. The iLF capability to support damage detection will be assessed by applying it to the four other cases (cases #2-5) and determining if a relationship exists between the absolute difference of the damaged and baseline $\omega_{n}$. In addition, the $\bm{\phi}_{n}$ change across the different cases will be used for damage localisation.
@@
Further, the COMAC is employed in Figure 7 to validate its applicability to damage localisation. A similar situation to that described in Figures 4 and 5 is observed: the COMAC values decrease significantly around nodes #2 and 4, i.e., the damaged area. In particular, for case #5, the COMAC deviation precisely detects the added mass location at node #3. Note that only the results for y-direction modal displacements are presented, for the sake of brevity, since a similar result is found for the z-direction.
Figure 7: Numerical case study: COMAC values between the iLF identified $\bm{\phi}_{n}$ of the baseline case and cases #2-5. (Figure 7(a)) Across the y-axis. (Figure 7(b)) Across the y-axis – Detail for COMAC $\in$ [0.995, 1].
@@ 4 Experimental case study
The BAE Systems Hawk T1A experimental dataset has recently been made available by the Dynamics Research Group at the University of Sheffield [41]. This dataset extends a previous database [64] that focused on the starboard wing of the same aircraft. The decommissioned full trainer jet aircraft, once in service with the British Royal Air Force, is shown in Figure 8 at the LVV facilities at the University of Sheffield. The data used in this study refer to the entire Hawk airframe structure, retrieved from [42]. The same dataset was also used previously to validate the MIMO extension of the LF for linear SI [43] and in [65] for the preliminary validation of Convolutional Neural Network with Reversed Mapping approach to SHM. Here in this work, the dataset is used to validate the proposed damage assessment procedure. For the sake of brevity, only the key aspects of the experimental setup, testing procedures, and recorded data used in this research work will be detailed. A more detailed discussion can be found in [41, 64].
Figure 8: BAE Systems Hawk T1A aircraft experimental setup at the LVV at the University of Sheffield (Retrieved from [41]).
@@ 4.1 Experimental setup
The airframe was equipped with 85 PCB Piezotronics (uniaxial and triaxial) accelerometers, featureing different nominal sensitivities ranging from 10 mVg-1 to 100 mVg-1, along with other sensors, totalling 139 recording channels. The sensor layout was designed to cover specific target components: the two wings, the two horizontal stabilisers, the vertical stabiliser, key points on the fuselage, and the three landing gears. For reasons described later, only the accelerometers located on the port wing have been considered in this study. None of the other sensors -– fibre-Bragg grating strain gauges, temperature sensors, etc. -– were considered. For MIMO linear SI, five TiraTM TV 51140-MOSP modal shakers [^2] https://www.tira-gmbh.de/fileadmin/inhalte/download/schwingprueftechnik/schwingpruefanlagen/modal/EN/100N_bis_2700N/data_sheet_system_tv_51140-MOSP_eng_V05.pdf were used, attached to PCB PiezotronicsTM 208C02 force transducers (sensitivity: 11.11 mVN-1) [^3] https://www.pcb.com/products?m=208c02, applying input forces to the respective locations as indicated in Figure 9.
Figure 9: BAE Systems Hawk T1A aircraft: Shaker positions for the MIMO tests on the starboard (Figure 9(a)) and the port wing ((Figure 9(b))), the horizontal stabiliser ((Figure 9(c))), and the vertical stabiliser (Figure 9(d)). Adapted from [41].
@@ 4.1.1 Testing procedures
The dataset contains white Gaussian noise (WGN) inputs, with five amplitude levels increasing from 0.1 V to 0.5 V. Pink Gaussian noise excitations are also available for linear SI. In contrast, odd random-phase multi-sine excitations can be used to investigate the nonlinear response of the system (which is out of the scope of this work). SIMO and MIMO tests were performed, roving the input source location for SIMO tests. However, these have not been considered here either since the LF was already proved viable for SIMO vibration-based dynamic monitoring in previous works [43].
@@ 4.1.2 Recorded data and selected recordings
    Six cases of added mass were considered. Specifically, these arise from adding different weights at three locations: the tip of the leading edge (PW_TLE), the root of the same (PW_RLE), and the centre of the trailing edge (PW_CTE), with the mass always added on the upper side, as shown in Figure 10. The added weights considered are M1 (254.3 g), M2 (616.8 g), and M3 (916.8 g).
@@
    The interest was focused on damage cases simulated on the port wing, which covered several types of damage. These include removed panels (at different locations) and added masses (located at different sites and with different weights to simulate growing damage severity).
@@
    All the relevant damage scenarios were separately tested.
@@
[…]
@@ 12 PR_WN15 PW5 panel removal
Moreover, aiming at MIMO SI and vibration-based SHM, these signals have all been compared to the health baseline for the same input amplitude (HS_WN5), using the 21 uniaxial 10 mVg-1 accelerometers included on the wing port. These sensors were recording in the vertical direction only, arranged in a grid of two lines on the top side (running parallel to the leading and trailing edges) and one line on the bottom side (parallel to the leading edge), with respectively of 14 and 7 uniaxial accelerometers, for a total of 21 output channels (see Figure 11).
@@ 4.2 Results
Before addressing the effectiveness of the proposed methodology for SHM purposes, it is worthwhile to investigate the accuracy of modal identification via iLF compared to that obtained with the benchmark methods SSI-CVA and LSCE on this experimental dataset. A further source of comparison is identified in the literature, i.e. the results reported in [43]. However, please consider that those results were obtained for the case with 0.4 V input voltage, while all data in this work refer to the 0.5 V input case. The first step involves identifying the stable modes for the healthy case, case #1, via iLF and benchmark methods. In Figure 12, the three stabilisation diagrams are presented. It is clear that the iLF (Figure 12(a)) identification succeeds in finding the most stable modes, while SSI-CVA (Figure 12(c)) and especially LSCE (Figure 12(b)) identifications return fewer and less stable poles.
Figure 12: BAE Systems Hawk T1A aircraft: Stabilisation diagrams for the modes identified via iLF (Figure 12(a)), LSCE (Figure 12(a)), and SSI-CVA (Figure 12(c)) for the baseline scenario.
From [43] and the current iLF identification, 41 modes were identified. However, only the first three are used here for the SHM task. Hence, in Table 3, the first three pairs of natural frequencies ($\omega_{n}$) and damping ratios ($\bm{\phi}_{n}$) identified in this work for the undamaged Hawk T1A Aircraft are presented alongside the aforementioned literature results in Table 3, while the iLF-identified $\omega_{n}$ and $\bm{\phi}_{n}$ for all cases are shown Figures 13(a) and 13(b).
Table 3: BAE Systems Hawk T1A aircraft: $\omega_{1-3}$ and $\zeta_{1-3}$ identified via iLF, SSI-CVA, and LSCE. 0.4 V input case refers to the literature result in [43] for a healthy case with a lower input amplitude (transposed).
Method Mode 1 Mode 2 Mode 3
Natural Frequency [Hz] – (Difference [%])
0.4 V input [43] 6.98 15.43 16.32
iLF 7.03 15.42 16.28
(0.7) (-0.1) (-0.3)
SSI-CVA
LSCE 16.25
(–) (–) (-0.4)
Damping Ratio [-] – (Difference [%])
0.4 V input [43] 0.03 0.01 0.01
iLF 0.03 0.01 0.01
(0) (0) (0)
SSI-CVA
LSCE 0.03
(–) (–) (200)
Figure 13: BAE Systems Hawk T1A aircraft: $\omega_{1-3}$ (Figure 13(a)) and $\zeta_{1-3}$ (Figure 13(b)) identified via iLF for the 12 cases under scrutiny.
@@
  1. Mode #1 1st bending (downwards);
  2. Mode #2 2nd bending (downwards);
  3. Mode #3 1st coupled between bending (downwards) and torsion.

### v4

@@ 2.3 A novel two-step modal & damage identification method
As many SHM approaches rely on identified modal parameters, a more accurate and efficient method can be developed by combining the two tools introduced earlier: the iLF and the MTMAC.
The proposed SHM workflow consists of:
@@
    Modal identification of the system;
@@
    Using the identified parameters to validate the MTMAC for damage assessment and quantification;
@@
    Applying the COMAC, for completeness, to determine the location of the damage.
To benchmark this method, the conventional LSCE and SSI-CVA techniques are employed, and their results are compared with those obtained using the iLF. In addition, a traditional vibration-based damage assessment, quantification, and localisation approach, based on comparing $\omega_{n}$ and $\bm{\phi}_{n}$, is applied to further validate the MTMAC as a damage index.
@@ 3 Numerical case study
To assess the suitability of the iLF and MTMAC for damage detection, a numerical case study is presented using a 2D hollow rectangular cross-section cantilever beam. The rationale here is to isolate the performance of the modal extraction algorithm (iLF) and damage index (MTMAC) in an intentionally simple benchmark, fully controlled, without unmodeled nonlinearities (e.g., joints and contacts) that can introduce confounding effects when interpreting damage indices. The paired full-scale aircraft experiment, which will be discussed in the next Section, will provide the necessary complexity for real-life validity.
The beam hollow rectangular cross-section is shown in Figure 1. It is made of aluminium, with a Young’s modulus (E) of 69 GPa and density ($\rho$) of 2700 kgm-3, and a length (L) of 1.8 m. The beam is discretised, in MATLAB, into four 2D (y and z displacements) Euler-Bernoulli beam elements of equal length. Sample beam element mass and stiffness matrices are available in Appendix A. The beam is fully constrained in all DoFs at the wall, $N_{1}$.
Figure 1: Numerical case study: schematic of the 2D beam with the dimensions of its hollow rectangular cross-section. Not in scale.
For the numerical MIMO experiment, two 1 N step input forces are applied in the y and z directions at node #2 ($N_{2}$$F_{y}$ and $F_{z}$ act at L/4 from the origin) at time zero of a 60 s recording. These and the displacements in the two directions are recorded at a sampling frequency of fs = 16384 Hz [^2] This frequency was required to simulate the resulting mass-spring-damper system using the lsim function (https://uk.mathworks.com/help/control/ref/dynamicsystem.lsim.html) in MATLAB, thus allowing the inspection of all modes up to 8192 Hz as per the Nyquist criterion. This enables the identification of all system modes. For all the modes, a 2% modal damping ratio is superimposed, using the uncoupled damping assumption. In addition to the baseline case (case #1) described above, the stiffness of the second element (see Figure 1) is reduced by 5% (case #2), 10% (case #3), and 20% (case #4) to simulate different damage scenarios on the beam. In addition, a fifth case (case #5) is considered, consisting of a mass addition of 100 g (just under 20% of the beam mass) to node 3, simulating an underwing payload or an engine pod mount assembly. This case allows us to examine the method sensitivity in distinguishing between damage and mass changes. The analytical results are obtained by simple eigenanalysis of the system mass and stiffness matrices. The modal damping is considered when computing the damped natural frequencies. For the identification process, the output displacement time series resulting from the excitation above are recorded and used to obtain the compliance FRFs in the usual fashion: Elementwise division of the output spectra by the input spectra. This process is repeated for the two inputs, thus resulting, in MATLAB notation, in an m:n:p matrix, where m is the number of outputs, n is the number of inputs, and p is the number of frequency bins. In this case, there are eight displacement output channels, one per direction (y and z) at each node, and two force input channels. LSCE and iLF require the FRFs, while the output time series are used with SSI-CVA. Table 1 shows the $\omega_{n}$ and $\zeta_{n}$ identified analytically and by applying the iLF and two traditional methods, SSI-CVA [44] and LSCE [54], to the baseline case. Notably, only the displacement DoFs are considered in the analysis. The percentage differences with respect to the analytical values are also reported (in brackets). The modal parameters are obtained with the help of stabilisation diagrams, as shown in Figure 2, for model orders ranging from 24 to 50. All numerical system analytical, numerical, and identified results are obtained in MATLAB 2024b.
@@
The maximum error of the identified parameters ($\omega_{n}$ and $\zeta_{n}$) reported in Table 1 is 0%. However, LSCE fails to identify lower modes in higher-dimensional systems with closely spaced modes, such as modes 1-2, as well as modes 3-6. These results align with expectations, as this issue is well-documented in the literature [52, 34]. Conversely, Figure 2 clearly shows that the stabilisation diagrams from iLF (Figure 2(a)) and SSI-CVA (Figure 2(c)) can accurately identify the stable modes, while for LSCE (Figure 2(b)) this is troublesome, resulting in only ten (out of sixteen) modes identified as stable. The stabilisation diagrams in question are constructed by varying the system order from $k_{\min}=32$ to $k_{\max}=60$ with an increment $\Delta k=2$, yielding the extraction set $\mathbf{k}=\{\,k_{\min}:\Delta k:k_{\max}\,\}$. Hard criteria are imposed through the admissible damping range $\zeta_{n}\in[\zeta_{\min},\zeta_{\max}]=[0.005,\,0.03]$ and the natural frequency range $\omega_{n}\in[\omega_{\min},\omega_{\max}]=[0,\,3800]$. Soft criteria include frequency stabilisation $\Delta\omega_{\mathrm{stab}}=0.01$, damping stabilisation $\Delta\zeta_{\mathrm{stab}}=0.05$, and MAC stabilisation $\mathrm{MAC}_{\mathrm{stab}}=0.95$. Additionally, the number of MAC comparisons $N_{\mathrm{MAC}}=10$ are considered in the selection process. It should be noted, although not included in Table 1, that the MAC between the analytical $\bm{\phi}_{n}$ and those identified via iLF and SSI-CVA is equal to, or very close to, 1. On the other hand, the LSCE identified $\bm{\phi}_{n}$ MAC, when compared to the analytical values, never exceeds 0.58.
@@ 3.1 Investigation of damage effects
Given the good accuracy and precision shown for the baseline case by the iLF when compared to the SSI-CVA and LSCE methods, only the iLF-derived modal parameters are used for damage assessment in the numerical case study. It must be noted that performing these using the analytical modal parameters yields the same results as those presented below, since the iLF identified modal parameters perfectly match their analytical counterparts. The iLF capability to support damage detection will be assessed by applying it to the four other cases (cases #2-5) and determining if a relationship exists between the absolute difference of the damaged and baseline $\omega_{n}$. In addition, the $\bm{\phi}_{n}$ change across the different cases will be used for damage localisation.
@@
In this sense, a needed remark concerns thresholding. The current study is designed as a proof-of-concept validation of the iLF and MTMAC framework, focusing on sensitivity and consistency across different (numerical and experimental) datasets. However, decision thresholds are natural in applied SHM. Thus, while this work is exploratory/validation-oriented, and the Authors do not want to overclaim, it is still possible to define some quantitative MTMAC values that could serve as illustrative thresholds (with the caveat that they are case-study dependent). More specifically, based on the outcomes shown in Figure 6 and discussed previously:
  • 1.
    Case #2 (5% stiffness reduction): MTMAC $\approx$ 0.05;
  • 2.
    Case #3 (10% stiffness reduction): MTMAC $\approx$ 0.102;
  • 3.
    Case #4 (20% stiffness reduction): MTMAC $\approx$ 0.208;
  • 4.
    Case #5 (mass addition equal to 2% of the beam whole mass, no stiffness damage): MTMAC $\approx$ 0.782.
It is possible to say, at least for the beam example:
  • 1.
    MTMAC $\leq 0.1-0.15$ may correspond to mild/local damage;
  • 2.
    MTMAC $\geq 0.15-0.2$ may correspond to clear/severe stiffness reduction ($\geq 20\%$);
  • 3.
    MTMAC $\gg 0.5$ may indicate very severe damage and/or a substantial structural change (e.g., a large added mass).
Finally, the COMAC is employed in Figure 7 to validate its applicability to damage localisation. A similar situation to that described in Figures 4 and 5 is observed: the COMAC values decrease significantly around nodes #2 and 4, i.e., the damaged area. In particular, for case #5, the COMAC deviation precisely detects the added mass location at node #3. Note that only the results for y-direction modal displacements are presented, for the sake of brevity, since a similar result is found for the z-direction.
Figure 7: Numerical case study: COMAC values between the iLF identified $\bm{\phi}_{n}$ of the baseline case and cases #2-5; (a) across the y-axis and (b) across the y-axis – Detail for COMAC $\in$ [0.995, 1].
@@ 4 Experimental case study
The BAE Systems Hawk T1A experimental dataset has recently been made available by the Dynamics Research Group at the University of Sheffield [41]. This dataset extends a previous database [75] that focused on the starboard wing of the same aircraft. The decommissioned full jet trainer aircraft, once in service with the British Royal Air Force, is shown in Figure 8 at the LVV facilities at the University of Sheffield, along with the global coordinate reference frame and its origin. The data used in this study refer to the entire Hawk airframe structure, retrieved from [42]. The same dataset was also used previously to validate the MIMO extension of the LF for linear SI [52] and in [76] for the preliminary validation of a Convolutional Neural Network with a Reversed Mapping approach to SHM. Here in this work, the dataset is used to validate the proposed damage assessment procedure. For the sake of brevity, only the key aspects of the experimental setup, testing procedures, and recorded data used in this research work will be detailed. A more detailed discussion can be found in [41, 75].
Importantly, this laboratory dataset includes multiple substructures, connections, and operational complexities typical of aeronautical assemblies, complementing the controlled numerical benchmark and addressing realism and complexity. In this context, it must be recalled that the System Identification approach used here (iLF) is intended for LTI systems; thus, structural nonlinearities cannot be explicitly modelled in this analytical formulation. However, this limitation is inherent in any modal-based SHM approach, as the Fourier Transform itself is a linear operator. Hence, nonlinearities in the mechanical system will manifest as noise-like distortions in the Fourier spectra of the signal and therefore in the Frequency Response Functions (FRFs) from which the modal parameters are estimated. For an overview of nonlinear modal analysis approaches to full-scale aeronautical systems, the interested reader can refer to [77, 78] for an application to a General Dynamics F-16. Nevertheless, the specific signals employed here correspond to the system vibrational response under low-amplitude broadband excitations, where the aircraft is proven to behave mostly in the linear regime (as a standard practice in ground vibration tests).
Figure 8: BAE Systems Hawk T1A aircraft experimental setup at the LVV at the University of Sheffield, showing the reference frame and its origin (adapted from [41]).
@@ 4.1 Experimental setup
The airframe was equipped with 85 PCB Piezotronics (uniaxial and triaxial) accelerometers, featuring different nominal sensitivities ranging from 10 mVg-1 to 100 mVg-1, along with other sensors, totalling 139 recording channels. The sensor layout was designed to cover specific target components: the two wings, the two horizontal stabilisers, the vertical stabiliser, key points on the fuselage, and the three landing gears. For reasons described later, only the accelerometers located on the port wing have been considered in this study. None of the other sensors -– fibre-Bragg grating strain gauges, temperature sensors, etc. -– were considered. For MIMO linear SI, five TiraTM TV 51140-MOSP modal shakers [^3] https://www.tira-gmbh.de/fileadmin/inhalte/download/schwingprueftechnik/schwingpruefanlagen/modal/EN/100N_bis_2700N/data_sheet_system_tv_51140-MOSP_eng_V05.pdf were used, attached to PCB PiezotronicsTM 208C02 force transducers (sensitivity: 11.11 mVN-1) [^4] https://www.pcb.com/products?m=208c02, applying input forces to the respective locations as indicated in Figure 9.
Figure 9: BAE Systems Hawk T1A aircraft: Shaker positions for the MIMO tests on the (a) starboard and the (b) port wing, the (c) horizontal stabiliser, and the (d) vertical stabiliser (adapted from [41]).
@@ 4.1.1 Testing procedures
The dataset contains white Gaussian noise (WGN) inputs, with five amplitude levels increasing from 0.1 V to 0.5 V. Pink Gaussian noise excitations are also available for linear SI. In contrast, odd random-phase multi-sine excitations can be used to investigate the nonlinear response of the system (which is out of the scope of this work). SIMO and MIMO tests were performed, varying the input source location for SIMO tests. However, these have not been considered here either since the LF was already proven viable for SIMO vibration-based dynamic monitoring in previous works [52]. The excitation amplitudes used here (0.5 V) ensure that the structure remains in its linear regime, as is standard in ground vibration testing.
@@ 4.1.2 Recorded data and selected recordings
    Six cases of added mass were considered. Specifically, these arise from adding different weights at three locations: the tip of the leading edge (PW_TLE), the root of the same (PW_RLE), and the centre of the trailing edge (PW_CTE), with the mass always added on the upper side, as shown in Figure 10. The added weights considered are M1 (254.3 g), M2 (616.8 g), and M3 (916.8 g);
@@
    The interest was focused on damage cases simulated on the port wing, which covered several types of damage. These include removed panels (at different locations) and added masses (located at different sites and with different weights to simulate growing damage severity);
@@
    All the relevant damage scenarios were separately tested;
@@
Figure 10: BAE Systems Hawk T1A aircraft: Simulated damage (PW_xxE) and removed panels (PWnn) positions on the port wing (PW) (adapted from [41]).
@@ 12 PR_WN15 PW5 panel removal
Moreover, aiming at MIMO SI and vibration-based SHM, these signals have all been compared to the health baseline for the same input amplitude (HS_WN5), using the 21 uniaxial 10 mVg-1 accelerometers included on the wing port. These sensors were recording in the vertical direction only, arranged in a grid of two lines on the top side (running parallel to the leading and trailing edges) and one line on the bottom side (parallel to the leading edge), with respectively 14 and 7 uniaxial accelerometers, for a total of 21 output channels (see Figure 11).
@@ 4.2 Results
Before addressing the effectiveness of the proposed methodology for SHM purposes, it is worthwhile to investigate the accuracy of modal identification via iLF compared to that obtained with the benchmark methods SSI-CVA and LSCE on this experimental dataset. A further source of comparison is identified in the literature, i.e. the results reported in [52]. However, please consider that those results were obtained for the case with 0.4 V input voltage, while all data in this work refer to the 0.5 V input case. The first step involves identifying the stable modes for the healthy case, case #1, via iLF and benchmark methods. In Figure 12, the three stabilisation diagrams are presented. It is clear that the iLF (Figure 12(a)) identification succeeds in finding the most stable modes, while SSI-CVA (Figure 12(c)) and especially LSCE (Figure 12(b)) identifications return fewer and less stable poles. The stabilisation diagrams are constructed by varying $k$ from $k_{\min}=84$ to $k_{\max}=150$ with $\Delta k=2$, yielding the extraction set $\mathbf{k}=\{\,k_{\min}:\Delta k:k_{\max}\,\}$. Hard criteria are imposed through the admissible ranges $\zeta_{n}\in[0.005,\,0.1]$ and $\omega_{n}\in[5,\,160]$. Soft criteria include $\Delta\omega_{\mathrm{stab}}=0.05$, $\Delta\zeta_{n,\mathrm{stab}}=0.3$, and $\mathrm{MAC}_{\mathrm{stab}}=0.95$. Additionally, $N_{\mathrm{MAC}}=3$ are considered in the selection process.
Figure 12: BAE Systems Hawk T1A aircraft: Stabilisation diagrams for the modes identified via (a) iLF, (b) LSCE, and (c) SSI-CVA for the baseline scenario.
From [52] and the current iLF identification, 41 modes were identified. However, only the first three are used here for the SHM task. Hence, in Table 3, the first three pairs of natural frequencies ($\omega_{n}$) and damping ratios ($\bm{\phi}_{n}$) identified in this work for the undamaged Hawk T1A Aircraft are presented alongside the aforementioned literature results in Table 3. Notably, only the iLF is able to identify all modes of interest, so SSI-CVA and LSCE results are disregarded from now onwards. The iLF-identified $\omega_{n}$ and $\bm{\phi}_{n}$ for all cases are shown Figures 13(a) and 13(b). All the processing carried out in this work for obtaining the modal parameters of the BAE Systems Hawk T1A is done in MATLAB 2024b.
Table 3: BAE Systems Hawk T1A aircraft: $\omega_{1-3}$ and $\zeta_{1-3}$ identified via iLF, SSI-CVA, and LSCE. 0.4 V input case refers to the literature result in [52] for a healthy case with a lower input amplitude.
Mode # Natural Frequency [Hz] -(difference [%]) Damping Ratio [-] -(difference [%])
0.4 V input [52] iLF SSI-CVA LSCE 0.4 V input [52] iLF SSI-CVA LSCE
1 6.98 7.03 - - 0.03 0.03 - -
(0.7) (-) (-) (0) (-) (-)
2 15.43 15.42 - - 0.01 0.01 - -
(-0.1) (-) (-) (0) (-) (-)
3 16.32 16.28 - 16.25 0.01 0.01 - 0.03
(-0.3) (-) (-0.4) (0) (-) (200)
Figure 13: BAE Systems Hawk T1A aircraft: (a) $\omega_{1-3}$ and (b) $\zeta_{1-3}$ identified via iLF for the 12 cases under scrutiny.
@@
  1. Mode #1: 1st bending (downwards);
  2. Mode #2: 2nd bending (downwards);
  3. Mode #3: 1st coupled between bending (downwards) and torsion.